Tom's father is 3 times old as Tom. 4 years ago, he was 4 times older. How old
is Tom? show your equation.

Answer by user_dude2008(1862)

f=3t

f-4=4(t-4)

3t-4=4(t-4)

3t-4=4t-16

3t-4t=-16+4

-t=-12

t=12
f=36

Answer: father is 36, tom is 12